-3-(-7)=
What are the rules when parentheses are used? Drop the sign within? It’s been forever!

Found 3 solutions by ewatrrr, ikleyn, Boreal:
Answer by ewatrrr(24785)

 
Hi,
 -3-(-7) is actually -3 + (-1)(-7)    |A minus number times a minus number


-3-(-7)= -3 + 7 = 4  


Wish You the Best in your Studies.


Answer by ikleyn(52878)

.

The rule is SIMPLE:


        when you see the sign " - " before parentheses,  you may take it off,  together with parentheses,

        but then you  MUST  change the signs of the terms  (numbers)  inside these parentheses to opposite.


            -3 - (-7) = -3 + 7 = 4.     ANSWER


Solved, answered and explained.     And completed.



Answer by Boreal(15235)

When there is confusion if one does not use them.
Also use them when you might change the order of doing operations that appears to violate the rules. If you had 4+9/2, this would be 4+4.5 or 8.5. But you might really mean (4+9)/2, which is 6.5. x-1/x is not understandable without parentheses; we see many problems written like this and have to guess what the person wants.
Two minus signs together without parentheses may appear as one. When I view the above, I connect the two minus signs and then put a vertical line through to make it +.
To answer the question, a parenthesis is not required, but it does help clarify. Don't drop the sign within.